# rtboxstyle linecolor=black, backgroundcolor=lightlight-gray, leftmargin=0cm, rightmargin=0cm, skipabove=4pt, skipbelow=4pt, linewidth=0.4pt, innertopmargin=3pt, innerbottommargin=3pt, innerrightmargin=6pt, innerleftmargin=6pt, roundcorner=2

### v1

@@ Abstract
Provenance graph analysis plays a vital role in intrusion detection, particularly against Advanced Persistent Threats (APTs), by exposing complex attack patterns. While recent systems combine graph neural networks (GNNs) with natural language processing (NLP) to capture structural and semantic features, their effectiveness is limited by class imbalance in real-world data. To address this, we introduce ProvSyn, an automated framework that synthesizes provenance graphs through a three-phase pipeline: (1) heterogeneous graph structure synthesis with structural-semantic modeling, (2) rule-based topological refinement, and (3) context-aware textual attribute synthesis using large language models (LLMs). ProvSyn includes a comprehensive evaluation framework that integrates structural, textual, temporal, and embedding-based metrics, along with a semantic validation mechanism to assess the correctness of generated attack patterns and system behaviors. To demonstrate practical utility, we use the synthetic graphs to augment training datasets for downstream APT detection models. Experimental results show that ProvSyn produces high-fidelity graphs and improves detection performance through effective data augmentation.
@@ 1 Introduction
Data provenance [29], which involves recording the history of data modifications, access, and usage, serves as a critical tool in cybersecurity systems for tracing attacks [61]. However, as system scales expand, the volume of provenance data increases dramatically, posing challenges for data management and analysis. To address this, researchers have begun to represent the provenance of the data using graph structures and applying graph machine learning algorithms for the detection and tracking of intrusions [21, 62, 53]. This approach effectively captures the complexity of system activities and enables the detection of attacks that traditional methods struggle to identify, such as zero-day attacks and APT attacks.
The quality of provenance data sets largely determines the effectiveness of provenance-based intrusion detection systems (IDS) trained on them. However, there are very few datasets available, and even fewer datasets that can accurately describe benign and malicious activities in a system. One possible solution is to produce system logs captured from real-world scenarios. However, one of the key disadvantages of using such datasets is privacy concerns. Sharing such dataset across the community could result in leakage of critical privacy information, and hence it’s difficult for organizations to release such datasets. Another solution is to collect benign and malicious activities in experimental environments. For example, the DARPA Transparent Computing (TC) [2] program contributes two widely used datasets in the cybersecurity community, DARPA TC dataset and DARPA OpTC dataset. However, such datasets, while giving us insights into system activities, convey limitations, one of which being the class imbalance issue [6], where less common or underrepresented program behaviors are under-sampled. This imbalance hampers the generalization ability of machine learning models trained on such data, limiting their effectiveness in unseen environments [77, 9, 76, 6].
To mitigate the issue of data imbalance, a common approach is to synthesize high-fidelity provenance data that expands coverage of minority classes and patterns. Such data synthesis can be achieved through methods such as graph generation models or large language models (LLMs), yet each has critical limitations. On the one hand, most graph generation models focus on structural connectivity and lack support for generating rich textual attributes, which are essential for downstream intrusion detection tasks [12, 71, 37, 53]. On the other hand, although LLMs are strong at producing semantically rich log entries, their outputs often fail to preserve the underlying graph structure when transformed into provenance graphs [4, 1].
To overcome these challenges, we propose ProvSyn, a hybrid framework that integrates the structural modeling ability of graph generation models with the text generation ability of LLMs to jointly synthesize provenance graphs with both accurate topology and meaningful textual attributes. Specifically, ProvSyn is a three-phase synthesis framework comprising structural synthesis, topological refinement, and textual attribution. The initial phase constructs heterogeneous graph topology by generating nodes and edges with categorical attributes. Subsequent structural refinement applies domain-specific rules to eliminate unreasonable connections and isolated entities. In the third phase, an LLM is used to generate textual attributes of nodes. However, the implementation of our framework confronts three principal technical challenges:
  • Scalability to provenance graphs. Provenance graphs are heterogeneous graphs with more than 10,000 nodes and edges, whereas most heterogeneous graph generation models are designed for molecular graphs and are typically trained on graphs with fewer than 100 nodes [72, 17]. Effectively scaling graph generation models designed for other domains to large-scale provenance graphs thus remains a challenge.
  • Textual attributes synthesis. First, LLMs inherently struggle with graph structure comprehension [33], necessitating transformation of provenance graphs into LLM-interpretable formats. Second, due to lacking prior knowledge of typical provenance entity names, direct LLM prompting produces simplistic, low-diversity outputs.
  • Evaluation framework gap. The absence of a multidimensional evaluation framework hinders comprehensive assessment of synthesized graphs, particularly in verifying the semantic correctness of attack patterns and system behavior within generated provenance structures.
To address the first challenge, we adopt a domain-agnostic heterogeneous graph generation network [17] to jointly model node and edge labels along with structural connectivity. To make the training process scalable to large graphs, we adopt a restart-based random walk for subgraph sampling. Through a novel restart mechanism, this sampling algorithm escapes local regions and captures subgraphs that better reflect the overall graph structure.
To address the second challenge, we convert the graph structure into depth-first search (DFS) sequences to enhance the LLM’s understanding of the graph. In addition, we also designed two different masking strategies, including fully masking and partially masking, to construct a training set for fine-tuning the LLM, thereby providing it with knowledge of entity names in the provenance graph.
To address the third challenge, we design a comprehensive framework that evaluates fidelity from five perspectives: structural, textual, temporal, embedding, and semantic correctness. For structural fidelity, we adopt five Mean Discrepancy (MMD)-based metrics. For textual fidelity, we use standard NLP metrics, including BLEU, GLEU, and ROUGE. For temporal fidelity, we use sequence comparison metrics including Longest Common Subsequence (LCS) and Dynamic Time Warping (DTW). For embedding fidelity, we utilize various embedding methods, including DeepWalk and Doc2Vec. To accurately assess semantic correctness—the most challenging aspect—we adopt a model-based approach. Specifically, we sample positive and negative triples from real provenance graphs and train a classification model using contrastive learning.
In the experiments, we used our proposed evaluation framework to evaluate the provenance data synthesized by LLMs, including GPT-4o and Claude 3.7, as well as by ProvSyn. The results show that ProvSyn achieves higher fidelity than these strong LLMs. Moreover, by generating 100% novel and 100% diverse graphs, ProvSyn effectively mitigates the data imbalance issue in provenance datasets. When applied to an intrusion detection system, models trained on datasets augmented with ProvSyn-generated graphs exhibited reduced false positive rates during testing.
@@
  • We propose ProvSyn, a three-phase provenance graph synthesis framework that enables the construction of high-fidelity heterogeneous graphs with rich textual attributes.
  • We design a comprehensive multi-dimensional evaluation framework that evaluates the fidelity of synthesized provenance graphs, including: structural, textual, temporal, embedding, semantic correctness.
  • We mitigate the data imbalance problem in provenance datasets by synthesizing novel and diverse provenance graphs, thereby enhancing the generalizability of downstream intrusion detection models.
  • The proposed generation and evaluation frameworks can also be extended to other domains involving Heterogeneous Information Networks (HINs), such as social networks and citation networks.
@@ 2 Background and Motivation
2.1 Class imbalance in Provenance Graphs
@@
We statistically analyzed the distributions of entity type and event type across four commonly-used public provenance datasets (in Appendix A Figure 3). Among the datasets, Cadets, Theia, and Trace are collected from DARPA Engagement 3, and Nodlink [37] is simulated provenance data collected on an Ubuntu system using Sysdig. The distributions exhibit a typical long-tail pattern and reveal significant data imbalance. In both entity type and event type, a single category accounts for over 50% of instances, exceeding 70% in some cases. In contrast, most other categories contribute less than 20%, with many below 5%. This imbalance further reflects the uneven distribution of system behaviors and attack patterns in most provenance datasets.
The class imbalance problem poses a significant challenge for learning-based intrusion detection [21, 62, 74, 22, 39]. Previous studies in the deep learning community have revealed that class imbalance has a detrimental effect on a model’s classification performance as well as its generalization abilities [8, 34]. When deployed in open-world out-of-distribution (OOD) environments, detection models trained on imbalanced datasets may significantly exhibit performance degradation [67, 25]. Hence, a more balanced dataset is vital for deep learning-based provenance-based IDS (PIDS) to achieve optimal results. However, this problem is largely ignored by the community, as there are very few studies considering this issue.
2.2 Problems with Existing Methods
@@
To mitigate class imbalance, conventional techniques such as undersampling [44], oversampling [47], and the Synthetic Minority Over-Sampling Technique (SMOTE)  [9] have been proposed. However, these methods are not suitable for provenance graph data. Undersampling breaks the causality chain between system entities, while oversampling does not add variety to benign activities and can cause overfitting. SMOTE, which relies on linear interpolation in the feature space, cannot be applied to provenance graphs due to overlapping distributions of different classes in the feature space.
Another approach for augmenting provenance datasets is to leverage graph generation models to synthesize provenance graphs. However, most existing models focus primarily on generating the structural connectivity among nodes and edges, without supporting additional node or edge attributes [12, 71]. In contrast, provenance graphs contain rich textual attributes associated with nodes and edges, and the text attributes are crucial features for downstream intrusion detection systems [37, 53]. Yet, to date, no graph generation model inherently supports the joint generation of both graph structure and textual attributes.
@@
With the recent advances in large language models, their capabilities in natural language understanding and generation have been widely demonstrated [4, 1]. This motivates the exploration of the use of LLMs to synthesiz provenance data. Specifically, LLMs can be tasked with creating new provenance log entries. These generated logs contain rich texual and semantic information. However, when converting LLM-generated logs into graphs and evaluating them using graph-based metrics, their structural fidelity falls significantly short of graphs produced by heterogeneous graph generation models (see details in Section 5.2). This indicates that although current LLMs perform well in text synthesis, they lack a strong capability for synthesizing accurate graph topology and structure.
To address these limitations, we propose ProvSyn, a framework that integrates the complementary strengths of graph generation models for structural synthesis and large language models for text attribute generation.
3 Related Work
Provenance-based Intrusion Detection System Provenance-based detection methods can be divided into three main categories [74]: rule-based, statistics-based, and learning-based techniques. Rule-based methods [24, 46, 28] rely on predefined heuristics derived from known attack patterns to identify malicious activities. Statistics-based [23, 61, 45] approaches analyze deviations in provenance graph elements by measuring their statistical anomalies. Learning-based techniques employ deep learning models to capture either normal system behavior [21, 62] or malicious patterns [41, 74, 39], framing APT detection as either a classification [39] or anomaly detection task [74, 22]. These learning-based methods further branch into sequence-based approaches [41], which focus on execution workflows, and graph-based [21, 62, 74, 22, 39] methods, which use graph neural networks to model entity relationships and detect behavioral anomalies.
Network Traffic Augmentation To address class imbalance in network traffic datasets, one line of work applies sampling techniques. Jiang et al. [32] adopt one-side selection to remove noisy majority samples and apply SMOTE [9] to augment minority instances. Liu et al. [42] propose the Difficult Set Sampling Technique to improve sample representativeness. Another line of research explores synthetic network traffic generation using deep learning. Generative models, especially GANs, have been employed for various scenarios, including flow-based traffic [55], manipulated traffic [11], and social media traffic [54], addressing challenges in realistic traffic synthesis [5].
LLM-driven Synthetic Data Generation LLM-based synthetic data generation methods can be categorized into three main types: prompt engineering-based approaches [16, 20, 57, 43, 70, 38, 68], which utilize task descriptions, condition-value prompts, and in-context examples to steer generation; multi-step generation-based approaches [64, 15, 14, 63, 27], which break down complex tasks either at the sample or dataset level; and knowledge-enhanced approaches [50, 10, 30], which incorporate external knowledge such as knowledge graphs or web sources to improve the factual quality of outputs. The evaluation of synthetic data includes both direct metrics, which measure faithfulness [36] and diversity [73], and indirect metrics, which assess performance on downstream tasks through benchmark testing [58] and open-ended evaluation using human [26] or model-based [65] methods.
@@ 4 ProvSyn Design
We implement the generation of provenance graphs in three stages. In the first stage, we train a heterogeneous graph generation model using real provenance graph dataset to produce the graph structure. After that, we apply rule-based post-processing to remove invalid edges and isolated nodes. In the third stage, we fine-tune a large language model to label the nodes with appropriate names.
4.1 Problem Definition
A system audit log dataset can be represented as a provenance graph, where nodes correspond to system entities and directed edges represent system events. Formally, a provenance graph is a directed heterogeneous graph $G=(V,E)$, where each node $v_{i}\in V$ is associated with a type $t_{i}=f(v_{i})$ and a name $n_{i}=h(v_{i})$, and each edge $e_{i}\in E$ is associated with a type $l_{i}=g(e_{i})$. Each system event is represented as a labeled 5-tuple:
$e_{i}=(t_{i},n_{i},t_{j},n_{j},l_{i})$
where $t_{i}$ and $t_{j}$ denote the types of the source and target nodes, $n_{i}$ and $n_{j}$ denote their names, and $l_{i}$ denotes the edge type. ProvSyn is to learn a distribution over such provenance graphs from real system audit data, enabling the generation of synthetic graphs that preserve both the structural and semantic properties of the original dataset. The generated graphs are also heterogeneous, with diverse node and edge types, and each node is assigned a detailed name.
The objectives of ProvSyn are threefold: (1) to synthesize high-fidelity provenance graphs. This is our primary goal, as high-quality provenance data is essential for training reliable detection models; (2) to mitigate data imbalance in existing provenance graph datasets. ProvSyn is designed to learn the underlying distribution of real-world data and generate samples that represent underrepresented scenarios, thereby improving data diversity; and (3) to enhance the performance of downstream intrusion detection systems. By augmenting existing datasets with ProvSyn-generated graphs, the coverage of system behaviors can be expanded, leading to improved generalization and robustness of IDS models.
Figure 1: ProvSyn Architecture. First, a heterogeneous graph generation model is employed to generate the initial structure of a provenance graph. Subsequently, the topology of the graph is refined according to predefined rules. Finally, a large language model is used to synthesize the textual attributes of the nodes within the provenance graph.
4.2 Real-World Provenance Graph Construction
We begin by parsing each log entry to extract five key elements: source node type, source node name, destination node type, destination node name, and edge type. These elements are used to construct a directed provenance graph, where each node is assigned two attributes (type and name), and each edge is labeled with a type attribute. To avoid duplication, we maintain a hash table indexed by the MD5 hash of node names.
For log formats with simple key-value structures, we extract these elements directly using predefined keys. For formats with more complex attribute representations of entities and interactions, we apply regular expressions to extract the target elements. Extracting node names is particularly challenging; to improve accuracy, we define type-specific regular expressions tailored to each node type, ensuring that a valid name is extracted for every node.
4.3 Graph Structure Generation

### v2

# No Data? No Problem: Synthesizing Security Graphs for Better Intrusion Detection

@@ Abstract
Provenance graph analysis plays a vital role in intrusion detection, particularly against Advanced Persistent Threats (APTs), by exposing complex attack patterns. While recent systems combine graph neural networks (GNNs) with natural language processing (NLP) to capture structural and semantic features, their effectiveness is limited by class imbalance in real-world data. To address this, we introduce ProvSyn, a novel hybrid provenance graph synthesis framework, which comprises three components: (1) graph structure synthesis via heterogeneous graph generation models, (2) textual attribute synthesis via fine-tuned Large Language Models (LLMs), and (3) five-dimensional fidelity evaluation. Experiments on six benchmark datasets demonstrate that ProvSyn consistently produces higher-fidelity graphs across the five evaluation dimensions compared to four strong baselines. To further demonstrate the practical utility of ProvSyn, we utilize the synthesized graphs to augment training datasets for downstream APT detection models. The results show that ProvSyn effectively mitigates data imbalance, improving normalized entropy by up to 35%, and enhances the generalizability of downstream detection models, achieving an accuracy improvement of up to 38%.
@@ 1 Introduction
Data provenance [inam2023sok] records system activities for cybersecurity attack tracing [wang2020you]. Recent deep learning-based pids (pids[han2020unicorn, wang2022threatrace, zengy2022shadewatcher, rehman2024flash, li2023nodlink] effectively capture system complexity and detect sophisticated attacks like zero-day and apt (apt) attacks. These systems typically rely on unsupervised or self-supervised learning to model normal system behaviors. However, their effectiveness is constrained by training dataset quality and diversity [11023260]. These systems assume training data adequately represents diverse runtime configurations, workload patterns, and benign activities. However, this assumption is often violated in real-world scenarios where publicly available datasets exhibit significant limitations.
Datasets that used for training and evaluating learning-based pids often suffer from data imbalance, where certain system behaviors are overrepresented while others are severely underrepresented. These datasets simulate cyber attacks primarily from controlled experimental environments such as the DARPA TC and OpTC datasets [darpa-E3]. Prior studies show that benign activities in these datasets predominantly originate from automated workload generators, producing highly monotonous and predictable background behaviors [11023260]. Consequently, models trained on such data tend to overfit to repetitive patterns and fail to generalize to diverse, unpredictable activities in production environments. This data imbalance hampers the generalizability of machine learning-based pids, limiting their effectiveness in diverse real-world deployments [zhou2023improving, chawla2002smote, zhao2021graphsmote, anjum2021analyzing, 11023260].
To synthesize provenance graphs for data augmentation, our key insight is that provenance graphs encode two complementary types of information: structural information capturing relationships and interaction patterns among system entities, and semantic information describing entity attributes and identities. Both are critical for learning-based intrusion detection, as malicious activities are jointly determined by structural patterns (e.g., abnormal execution chains) and semantic attributes (e.g., suspicious file paths or command-line arguments) [li2023nodlink, rehman2024flash]. However, existing synthesis approaches fail to handle this dual nature. Graph synthesis methods [you2018graphrnn, goyal2020graphgen, de2018molgan] focus on generating topologies and categorical labels while neglecting rich textual attributes essential for intrusion detection. Conversely, LLM-based text generation excels at producing semantically rich log entries but fails to preserve graph structure when converted to provenance graphs, resulting in structurally invalid or semantically inconsistent connections. Therefore, effective synthesis must jointly model both structural connectivity and semantic attributes to generate high-fidelity provenance graphs for robust detection systems.
To address these limitations, we propose ProvSyn, a hybrid provenance graph synthesis framework integrating graph generation models and large language models. ProvSyn jointly models both structural and semantic information through a two-phase pipeline. In the graph structural synthesis phase, a heterogeneous graph generation model constructs topologies with accurate node and edge type distributions, capturing structural patterns from real provenance data. In the textual attribute synthesis phase, a fine-tuned LLM generates realistic textual attributes for each node, ensuring semantic consistency with the graph structure. However, realizing this framework introduces three principal technical challenges:
Scalability to provenance graphs. Provenance graphs often contain millions of nodes and edges, whereas most existing heterogeneous graph generation models are trained on graphs with fewer than 100 nodes. To address this scalability challenge, we adopt rwr (rwr) sampling to extract representative subgraphs. Unlike simple random walks that become trapped in densely connected regions, rwr periodically restarts from the initial node, enabling broader exploration. We then train GraphGen [goyal2020graphgen] on sampled subgraphs to jointly model node types, edge types, and structural connectivity. To guarantee validity and realism, we apply domain-specific refinement rules that eliminate invalid connections (e.g., process-clone-network).
Textual attributes synthesis. Although LLMs demonstrate strong text generation capabilities, they struggle with graph structure comprehension [jin2024large] and lack domain-specific provenance knowledge. Consequently, directly prompting LLMs often produces generic or unrealistic outputs. To address this, we first design a sft (sft) strategy adapting LLMs to the provenance domain. Then we convert provenance graphs into dfs (dfs) sequences, transforming graph structure into a linear format suitable for LLM processing while preserving connectivity and relationships.
Evaluation framework gap. A critical challenge in provenance graph synthesis is the absence of comprehensive evaluation frameworks. Existing methods [you2018graphrnn, li2018learning, goyal2020graphgen], focusing on topological properties are inadequate for provenance graphs requiring verification of both structural validity and semantic plausibility. To address this gap, we propose a novel five-dimensional fidelity evaluation framework, including structural, textual, temporal, embedding, and semantic metrics. This framework enables rigorous assessment of synthesized provenance graphs and addresses a critical gap in the graph generation literature.
In the experiments, we evaluate the fidelity of provenance data synthesized by ProvSyn following our proposed evaluation framework. We compare ProvSyn against four advanced LLMs (GPT-5.1 [openai-gpt51], Claude-Sonnet-4.5 [claude], DeepSeek-v3.2 [liu2025deepseek], and Qwen3-Max [yang2025qwen3]) on six widely used provenance datasets. The results show that ProvSyn achieves higher fidelity than directly using these LLMs from all five evaluation perspectives. Moreover, data imbalance is consistently mitigated after incorporating ProvSyn, improving normalized entropy by up to 35%. We further demonstrate the practical utility of ProvSyn by augmenting existing datasets with the synthesized graphs for improving four latest downstream APT detection models (NodLink [li2023nodlink], Magic [jia2024magic], Flash [rehman2024flash], Velox [bilot2025sometimes]). Experimental results demonstrate that detection models trained on ProvSyn-augmented datasets yield up to a 38% gain in accuracy, while concurrently bolstering discriminative reliability and diagnostic utility.
@@
  • We propose ProvSyn, a novel hybrid provenance graph synthesis framework that jointly models structural topology via graph generation models and textual semantics via fine-tuned LLMs, enabling the construction of high-fidelity provenance graphs.
  • We introduce a novel five-dimensional fidelity evaluation framework specifically designed for provenance graph synthesis, encompassing structural, textual , temporal, embedding, and semantic metrics, addressing a critical gap in existing graph generation.
  • Through systematic experiments on six benchmark datasets, we demonstrate that ProvSyn effectively mitigates data imbalance and enhances the generalizability of downstream apt detection models.
  • We open source ProvSyn at https://anonymous.4open.science/r/OpenProvSyn-4D0D/ to facilitate future research in provenance graph synthesis and pids.
@@ 2 Background and Motivation
2.1 Provenance Graph
Provenance graphs are constructed from system audit events that record interactions among system entities, such as processes, files, and network connections. Formally, a provenance graph is a directed heterogeneous graph with textual attributes $G=(V,E)$, where each node $v_{i}\in V$ is associated with a type $t_{i}$ and a textual attribute: a name $n_{i}$, and each edge $e_{i}\in E$ is associated with a type $l_{i}$. Each system audit event is represented as a labeled 5-tuple: $e_{i}=(t_{i},n_{i},t_{j},n_{j},l_{i})$, where $t_{i}$ and $t_{j}$ denote the types of the source and target nodes, $n_{i}$ and $n_{j}$ denote their names, and $l_{i}$ denotes the edge type. Edges are directed from the source node to the target node, reflecting the causal relationships between entities.
For node types, the most common node types include process, file, and network. For node names, process nodes are typically command lines, file nodes are their absolute file paths, and network nodes are their IP addresses. For edge types, common process–process interactions include CLONE, SENDMSG, and RECVMSG, which capture process creation and message-based communication. Common process–file interactions include READ, WRITE, and OPEN, which represent file access and modification behaviors. Common process–network interactions include CONNECT, SENDTO, and RECVFROM, which describe network connection establishment and data transmission.
2.2 Class Imbalance
Class imbalance poses a significant challenge in training effective machine learning models across diverse tasks, including financial fraud detection [dal2015calibrating, roy2024frauddiffuse], hate speech identification [davidson2017automated, caselli2021hatebert], and rare disease diagnosis [esteva2017dermatologist, tiu2022expert]. Class imbalance refers to a dataset characteristic where the number of samples across classes is uneven, often with one or more classes dominating the dataset while the remaining classes account for only a small fraction [japkowicz2002class]. Previous studies have revealed that class imbalance has a detrimental effect on a model’s classification performance as well as its generalization abilities in various machine learning tasks [buda2018systematic, johnson2019survey]. When deployed in open-world out-of-distribution (OOD) environments, detection models trained on imbalanced datasets may significantly exhibit performance degradation [yang2024generalized, he2009learning]. Hence, bridging the class imbalance gap in existing training datasets is vital for machine learning models to achieve robust results.
Figure 1: Entity Type and Event Type Distribution in Provenance Dataset in OPTC-H501.
@@ 2.2 Class Imbalance
To further illustrate the class imbalance issue in provenance datasets, we statistically analyze the distributions of entity types and event types across six commonly used public provenance datasets, including Cadets-E3, Theia-E3, Clearscope-E5, Theia-E5, OpTC-H201, and OpTC-H501. Figure 1 presents the distribution of OpTC-H501, while the distributions of the other datasets exhibit comparable distributional patterns and are shown in Appendix A (Figure 4). These distributions exhibit a typical long-tail pattern and reveal substantial data imbalance. For both entity types and event types, a single category accounts for more than 50% of the instances, exceeding 80% in some cases. In contrast, most other categories contribute less than 5%. This imbalance further reflects the uneven distribution of system behaviors in most provenance datasets.
2.3 Limitations of Existing Methods
@@
To mitigate class imbalance, conventional techniques such as undersampling [liu2008exploratory], oversampling [mohammed2020machine], and the Synthetic Minority Over-Sampling Technique (SMOTE) [chawla2002smote] have been widely adopted. However, these methods are unsuitable for provenance graphs. Undersampling discards majority-class samples, disrupting causality chains between system entities. Oversampling replicates minority-class instances without introducing diversity, potentially causing overfitting. SMOTE generates synthetic samples through linear interpolation between neighbors in feature space; however, due to overlapping class distributions, it cannot be effectively applied to provenance graphs where structural validity must be preserved.
An alternative approach leverages graph generation models to synthesize provenance graphs by learning underlying topology distributions. While frameworks such as MolGAN [de2018molgan], GraphRNN [you2018graph], and GraphGen [goyal2020graphgen] effectively capture structural connectivity, they primarily optimize adjacency patterns and categorical labels while neglecting textual attributes. In contrast, provenance graphs contain rich textual attributes associated with nodes—such as command lines, file paths, and IP addresses—that are crucial for downstream intrusion detection systems [li2023nodlink, rehman2024flash]. To date, no graph generation model inherently supports joint generation of both graph structure and textual attributes.
@@
Recent advances in LLMs demonstrate strong capabilities in natural language understanding and generation [achiam2023gpt, claude], motivating exploration of LLM-based provenance data synthesis. Specifically, LLMs can generate provenance log entries containing rich textual and semantic information. However, when converting LLM-generated logs into graphs and evaluating them using graph-based metrics, their structural fidelity falls significantly short of graphs produced by heterogeneous graph generation models (see Section 5.2). This indicates that although current LLMs perform well in text synthesis, they lack strong capability for synthesizing accurate graph topology and structure.
Figure 2: ProvSyn Architecture. First, a heterogeneous graph generation model constructs the initial structure, which is subsequently refined through rule-based topological constraints. Next, an LLM is fine-tuned on serialized provenance data to capture domain-specific semantics, enabling it to populate the generated nodes with synthesized textual attributes. Finally, the resulting graphs are validated via a fidelity evaluation framework across five distinct dimensions.
3 Threat Model and Design Goal
Threat Model. We follow the same threat model as past pids work [li2023nodlink, rehman2024flash, jia2024magic, jiang2025orthrus]. We assume attackers originate externally and aim to compromise internal systems for valuable data or persistent control. Their actions, though potentially complex, are assumed to leave observable traces in system logs monitored by kernel-level tools. The system hardware, OS, and core auditing software form our trusted base. Poison attacks and evasion attacks are not considered in our threat model.
Design Goal. ProvSyn is designed as a framework that integrates the complementary strengths of graph generation models for structural synthesis and large language models for text attribute generation. The objectives of ProvSyn are threefold: (1) to synthesize high-fidelity provenance graphs. High-fidelity provenance graphs require accurate modeling of both structural patterns and texual attributes; (2) to mitigate data imbalance in existing provenance graph datasets. ProvSyn is designed to learn the underlying distribution of real-world data and generate samples that represent underrepresented scenarios, thereby improving data diversity; and (3) to enhance the performance of downstream pids. By augmenting existing datasets with ProvSyn-generated graphs, the coverage of system behaviors can be expanded, leading to improved generalization and robustness of pids models.
@@ 4 ProvSynDesign
In this section, we introduce the design details of ProvSyn. Figure 2 shows the workflow of ProvSyn, which consists of three main components: graph structure synthesis, textual attribute synthesis, and fidelity evaluation framework. In the first stage, we train a heterogeneous graph generation model using real provenance graph dataset to generate the graph structure. In the second stage, we fine-tune a large language model to label the nodes with appropriate text attributes. Finally, we evaluate the fidelity of the synthesized provenance graphs from five perspectives.
4.1 Graph Structure Synthesis
In the first stage, we focus on generating the structure of provenance graphs. We first train a graph generation model on real provenance graph datasets, introduing the subgraph sampling strategy to address challenge of the scalability issue of large provenance graphs. Then for the synthesized graphs, we apply domain-specific refinement rules to ensure the validity of the generated graph structures.